# Electron-Phonon interaction and lattice thermal conductivity from metals to 2D Dirac crystals: a review

### v1

@@ 6 Conclusion
Across the metallic regime, first-principles calculations revise the long-held view that lattice heat flow is negligible: once e-ph scattering is included, phonons can account for up to $\sim\!40\%$ of $\kappa_{\mathrm{tot}}$, with the magnitude set primarily by $N(E_{\mathrm{F}})$, deformation potentials, and phonon frequencies. In semiconductors, solving the electron and phonon BTEs on equal footing resolves phonon drag, electron drag, and ultrafast non-equilibria, bringing theory into quantitative agreement with TDTR and isotope-controlled benchmarks (e.g., doped Si, GaAs, and polar nitrides). In 2D Dirac materials, symmetry, strain, carrier density, and finite size jointly establish a branch-selective hierarchy: in pristine suspended graphene, first-order ZA/ZO coupling is symmetry-forbidden, whereas substrates, edges, or tension activate/modify flexural channels and can shift the balance between in-plane and out-of-plane heat carriage. Moreover, when $E_{\mathrm{F}}<<k_{B}T$, four-particle processes become essential in 2D Dirac crystals, exposing a rich arena for engineering thermal rectifiers through controlled symmetry breaking and higher-order coupling. This opens a rich arena for engineering thermal rectification via controlled symmetry breaking and tailored higher-order e-ph coupling.
Looking forward, the path is clear: enlarge ab initio datasets to include correlated and magnetic Dirac sheets; embed fully non-local, frequency-dependent screening in van der Waals heterostructures so that strong e–ph coupling is treated beyond the Thomas–Fermi limit; incorporate second-order (four-particle) scattering—including EP $\leftrightarrow$ EP channels—into coupled electron- and phonon-BTE solvers; and couple these strong-coupling formalisms to femtosecond pump–probe experiments for real-time validation. Meeting these challenges will transform electron–phonon theory from a diagnostic tool into a genuinely predictive, strong-interaction design platform—capable of atomically precise control of heat and charge in next-generation electronic, photonic and quantum devices.
@@ Acknowledgments
We thank the Perimeter Institute for Theoretical Physics for providing a stimulating research environment in which this work was carried out. We acknowledge the Anishinaabek, Haudenosaunee, L¯unaap´eewak and Attawandaron peoples, on whose traditional lands Western University is located.

### v2

@@ 6 Conclusion
Across the metallic regime, first-principles calculations revise the long-held view that lattice heat flow is negligible: once e-ph scattering is included, phonons can account for up to $\sim\!40\%$ of $\kappa_{\mathrm{tot}}$, with the magnitude set primarily by $N(E_{\mathrm{F}})$, deformation potentials, and phonon frequencies. In semiconductors, solving the electron and phonon BTEs on equal footing resolves phonon drag, electron drag, and ultrafast non-equilibria, bringing theory into quantitative agreement with TDTR and isotope-controlled benchmarks (e.g., doped Si, GaAs, and polar nitrides). In 2D Dirac materials, symmetry, strain, carrier density, and finite size jointly establish a branch-selective hierarchy: in pristine suspended graphene, first-order ZA/ZO coupling is symmetry-forbidden, whereas substrates, edges, or tension activate/modify flexural channels and can shift the balance between in-plane and out-of-plane heat carriage. Moreover, when $E_{\mathrm{F}}<<k_{B}T$, higher-order processes become essential in 2D Dirac crystals, exposing a rich arena for engineering thermal rectifiers through controlled symmetry breaking and higher-order coupling. This opens a rich arena for engineering thermal rectification via controlled symmetry breaking and tailored higher-order e-ph coupling. Table 2 synthesizes, across the material classes treated here, the dominant phonon scattering mechanisms and the key control parameters governing lattice thermal conductivity at $T\simeq 300$$500\,\mathrm{K}$.
Table 2: A synthesized summary of dominant phonon scattering channels and controlling levers at $T\!\approx\!300$500500 K. Entries reflect trends compiled in this review (see Secs. 3–5).
Material class Dominant channels at 300–500 K Why (key levers / modeling notes)
Metals (noble $\to$ $d$-band; intermetallics) e–ph sets leading phonon lifetimes; 3ph anharmonicity captures intrinsic lattice resistance; $\kappa_{\mathrm{ph}}$ share: often $10\%$ (nobles), can reach $\sim\!10$$40\%$ in several $d$-band/intermetallic systems (once e–ph is included) High $N(E_{F})$ and screening shorten phonon MFPs ($\sim$ few–tens nm); optical branches in multi-atom bases add channels. Notes: Boundary/size effects matter for thin films.
Semiconductors (covalent vs polar) 3ph Umklapp dominates $\kappa_{\mathrm{ph}}$ (LA/TA in covalent; plus polar LO–TO coupling effects on phonon scattering in polar crystals); isotope and boundary scattering can be non-negligible. e–ph damping of phonons becomes relevant at high carrier densities Bonding character (short- vs long-range), carrier density, dielectric environment (screening). Notes: At 300 K, coupled e–ph BTE mainly impacts Seebeck/drag rather than $\kappa_{\mathrm{ph}}$ in many cases.
2D Dirac (graphene and kin) With mirror symmetry $\sigma_{h}$ intact, selection rules suppress ZA couplings, leading to LA/TA dominance; ZA is also suppressed in certain Dirac crystals such as graphene nanoribbons; substrates/gating/tension activate ZA channels and modify screening; nonadiabatic effects can arise at high $n$ or near flat-band limits Symmetry selection rules (ZA protection), environment-dependent screening (gate/encapsulation), strain/tension (ZA linearization). Notes: higher order in-plane scattering becomes important for 2D Dirac crystals at higher $T$; coherence (unified) term is usually negligible in simple monolayers.
Moiré/twisted graphene Enhanced e–ph in flat minibands; additional “moiré-Umklapp” channels; moiré phonons restructure phase space; coherences may contribute when inter-branch splittings $\sim$ linewidths Flat-band DOS and modified screening; moiré-phonon mini-bands; remote interfacial phonons (substrates). Notes: Start with Peierls–BTE; include coherence (unified) corrections when branches are quasi-degenerate; assess 4ph at high $T$.
[…]
@@ Acknowledgments
We thank the Perimeter Institute for Theoretical Physics for providing a stimulating research environment in which this work was carried out. We acknowledge the Anishinaabek, Haudenosaunee, Lūnaapéewak and Attawandaron peoples, on whose traditional lands Western University is located.